OpenStudy (anonymous):

Find each sum or difference. Simplify if possible. 10 1/8+3 3/4

OpenStudy (anonymous):



OpenStudy (fibonaccichick666):

So change the mixed fractions to improper fractions

OpenStudy (anonymous):

can you explain

OpenStudy (fibonaccichick666):

can you try, you didn't on the last one

OpenStudy (anonymous):

13 4/12

OpenStudy (fibonaccichick666):

so you need to change just \(10\frac{1}{8}\) into an improper fraction, not combine everything(ps, the lcm is not 12 here).

OpenStudy (fibonaccichick666):

So you need to figure out \[10\frac{1}{8}=\frac{?}{8}\]

OpenStudy (anonymous):

i dont know

OpenStudy (alexandervonhumboldt2):

do you still need help?